Circular No. 6404

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

CYGNUS X-1 AND GRS 1915+105 W. Cui, Center for Space Research, Massachusetts Institute of Tech-

nology, reports for the RXTE/ASM team: “Observations with the RXTE All Sky Monitor (ASM) reveal that Cygnus X-1 is in a bright flaring state. It was measured at ∼ 0.5 Crab (2–12 keV) before it started to brighten on May 10. The intensity reached ∼ 1 Crab on May 17 and ∼ 2 Crab on May 19. Another blackhole candidate, GRS 1915+105, has also become very bright. Early this year, this source was in a highly variable state (IAUC 6392). Its peak-to-peak amplitude became about a factor of 2 smaller between Apr. 7 and May 17, and its mean brightness decreased linearly. It was measured at ∼ 0.8 Crab on May 17 and has since brightened, reaching ∼ 3 Crab on May 19.”

COMET C/1990 N1 (TSUCHIYA-KIUCHI) K. Dennerl, J. Englhauser, and J. Trümper, Max-Planck-Institut für

Extraterrestrische Physik, Garching, communicate: “X-ray emission from comet C/1990 N1 (= 1990i = 1990 XVII) was detected in data obtained with the Positional Sensitive Proportional Counter (PSPC) of ROSAT during the all-sky survey. Between 1990 Nov. 18.76 and 20.36 UT (when Δ = 1.08 AU, r = 1.38 AU, and visual m_{1} ∼ 7.0), the PSPC was scanning across the comet, yielding a total exposure of 500 s. The x-ray image, corrected for the comet’s motion, shows an extended source at the expected position, with a radial extent of at least 10′, or 470 000 km. This PSPC observation allows, for the first time, spectroscopic studies of the x-ray radiation of a comet. The spectrum can be modelled by thermal bremsstrahlung emission with kT = 0.4 ± 0.1 keV. Inclusion of the 525-eV oxygen-fluorescence line improves the confidence of the fit. Pure line emission due to oxygen and carbon fluorescence can be ruled out. The 0.1- to 2.4-keV x-ray luminosity was 2.5 × 10^{16} erg/s and thus comparable to that of comet C/1996 B2 (cf. IAUC 6373). The optical luminosity of C/1990 N1, however, was lower by a factor of 15 than that of C/1996 B2. This second x-ray detection of a comet indicates that x-ray emission is a class property of such objects.”

COMET C/1996 E1 (NEAT) Visual m_{1} estimates by K. Sárneczky, Ráktanya, Hungary (0.44-m re-

flector): Apr. 20.84 UT, 13.4; May 18.88, 13.7.

1996 May 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green